# RS-Prune: Training-Free Data Pruning at High Ratios for Efficient Remote Sensing Diffusion Foundation Models

### v1

@@ 1 Introduction
In recent years, generation models, especially diffusion models [25, 13], have achieved remarkable progress in fields such as computer vision [27], medical imaging [35], and remote sensing (RS) [9]. Within the RS domain, generation models have been widely applied to data augmentation, image reconstruction, super-resolution, and high-resolution image synthesis, supporting practical applications in urban planning, land-use monitoring, and disaster response [4]. A powerful RS diffusion foundation model can provide a robust data and modeling backbone to further enhance these applications.
However, the effective training of RS diffusion foundation models critically depends on the quality and distribution of training data. The emergence of large-scale open-source datasets (e.g., Git-10M [20], RS5M [51]) provides valuable resources, yet it also introduces several critical challenges, including image redundancy, low-quality samples (e.g., noise and cloud cover), class imbalance [7], and scene homogeneity [40]. These issues not only hinder the training efficiency of RS foundation models but also limit the effectiveness of the resulting pretrained models on downstream tasks.
Recently, several studies in RS generation tasks have made preliminary explorations into data processing. For instance, RSDiff employs size cropping and noise augmentation [33]. WHU-RS19 [2] removes cloud and open-ocean regions to reduce low information content. Super-resolution works [15] often prioritize urban areas. Nevertheless, these methods mainly rely on rule-based or simplistic strategies to eliminate redundant data, without considering the specific dependence of generation models on data distribution [11] or the inherent characteristics of RS imagery, such as diversity, heterogeneity, and class balance. Moreover, many data pruning methods developed in the computer vision domain rely on scoring mechanisms [45] from supervised pretrained models [26] or retrieval of nearby images based on ImageNet benchmarks [24]. In contrast, a standardized labeled dataset for RS, comparable to ImageNet [8] in computer vision and spanning multiple resolutions and modalities, is currently lacking. Overall, systematic research on data selection for RS diffusion foundation models remains scarce.
@@
The effectiveness of diffusion foundation models depends more on the quality and distribution of the training data than on dataset size alone. Recent studies [6] have shown that diffusion models retain strong generation performance even when a large fraction (e.g., 90% of ImageNet samples) of the training data is removed, highlighting that a significant portion of the data contributes minimally. In the RS domain, raw data often contains substantial redundancy and noise, meaning that merely increasing the dataset size does not guarantee proportional performance gains. If generative models are trained on large volumes of data without appropriate selection, low-quality and redundant samples can not only slow convergence and increase computational and time costs, but also introduce distribution shifts that ultimately degrade model performance.
@@ 2.1 Diffusion foundation Models
Diffusion foundation models [32] play a pivotal role in image generation, data synthesis, and image reconstruction. For instance, Stable Diffusion (SD) 1.5 [29] leverages latent-space representations together with a UNet backbone, serving as a lightweight generation foundation model that provides strong pretrained initialization for ControlNet [47] and related low-level vision tasks [30]. More recently, Transformer-based architectures such as DiT [25] have emerged, while models like SD3 [10] and Flux [3] offer substantially stronger generation capabilities and higher fidelity, further advancing the scalability and versatility of diffusion frameworks.
In the RS domain, diffusion foundation models have also started to gain attention. To address the unique properties of RS imagery, such as multi-spectral and multi-resolution observations, geospatial information, and global coverage, many works have explored pre-training or fine-tuning foundation models on RS datasets [38, 39, 43]. For example, DiffusionSat [16] introduced geolocation as conditioning information and fine-tuned Stable Diffusion 1.5 on multi-source RS data, supporting multiple RS generation tasks. Meanwhile, models such as SR3 [31] and CDM [14], demonstrated strong super-resolution performance on natural images, providing a foundation for their adaptation to RS. Despite these advances, existing RS generation models typically aggregate multiple classification datasets or apply simplistic deduplication [22], thereby overlooking the distributional requirements of generation foundation modeling as well as the inherent heterogeneity and diversity of RS imagery. Consequently, the field lacks systematic exploration of data pruning strategies specifically designed to address the characteristics of remote sensing data and the requirements of RS generative foundation models.
@@ 2.2 Data Pruning Methods
Training data are critical to constructing RS diffusion foundation models. However, existing studies [21, 33] in this field have adopted simplistic deduplication and preprocessing strategies, such as removing cloud and ocean regions [2] and prioritizing urban areas [48]. These approaches are insufficient to address the sensitivity of generation models to data distribution, heterogeneity, and class imbalance.
Current data pruning methods can be broadly categorized into three types. First, data-valuation methods assign an importance score to each sample and select samples accordingly. For example, MoSo [37] estimates the change in empirical risk when a sample is removed. Although such methods are generally efficient, their performance can be affected by group effects and may lack generalization in complex real-world settings. Second, distribution-based methods rely on the geometric structure of the dataset. For instance, Moderate-DS [42] selected samples near the median. CCS [52] balanced data distribution and sample importance during selection. Finally, optimization-based methods leverage optimization techniques to guide sample pruning, such as temporal dual-depth scoring [50], gradient matching [17], scalable self-supervised pruning metrics [36], influence functions [18], and bilevel optimization [5].
@@
Beyond explicit pruning algorithms, recent foundation models have also explored large-scale data pruning pipelines. For example, DINOv2 [24] constructs its curated LVD-142M dataset by performing PCA-hash and copy-detection based deduplication, followed by self-supervised retrieval based on ImageNet benchmarks. Similarly, V-JEPA 2 [1] adopts a retrieval and reweighting strategy to align large-scale video data with target distributions.
@@ 3.2 Stage I: Entropy-based Pruning
We measure the information value of each image by computing its global Shannon entropy [34]. Specifically, for an image II, we compute its grayscale entropy $H(I)$ to capture the diversity of pixel intensities:
@@ 3.3 Stage II: Scene-Aware Clustering and Sampling
We employ a unified feature extractor $f(\cdot)$, namely Git-RSCLIP [20], with $\ell_{2}$ normalization to embed all images from standard datasets. Git-RSCLIP is specifically pretrained on large-scale remote sensing corpora and achieves substantially better performance than previous RS-oriented CLIP variants such as RemoteCLIP and GeoRSCLIP.
@@ 4.1 Datasets and evaluation
Moderate-DS [42] 44.07 92.40 0.3952 132.85 0.6222 199.24 124.86 0.5098 233.13 0.6730
$\mathrm{Cluster}_{\text{nearest}}$ [6] 45.24 89.58 0.3898 133.18 0.6237 201.36 150.11 0.5870 256.40 0.6804
Dinov2 [24] 39.34 91.76 0.3930 137.55 0.6121 180.98 127.73 0.5472 238.80 0.6761
A2-MAE [48] 45.96 90.39 0.3938 130.64 0.6013 360.74 148.36 0.6028 256.39 0.6993
@@
To evaluate the effectiveness of our data pruning approach for RS generation tasks, we conduct experiments on both global-scale and urban-scale datasets. Specifically, we use two representative global-scale optical datasets (i.e., GIT-10M and RS5M) that were carefully curated for RS generative foundation models, as well as a large-scale urban-scale multispectral dataset (i.e., USA-1m) collected in this work based on U.S. urban boundary products [19]. The global datasets provide broad coverage and diversity, while the urban dataset offers finer spatial resolution and richer texture information. This design allows us to assess the effectiveness and generalizability of our data pruning strategy across datasets with varying resolutions, modalities, volumes, and sensors.
  • Git-10M [20] A global-scale dataset comprising 10.5M satellite images (RGB) with a spatial resolution of 0.5 to 128 m, which spans multiple continents and geographic regions, covering diverse land-cover types such as urban areas, croplands, forests, mountains, and deserts.
  • RS5M [51] A global-scale dataset initially collected from 11 publicly available image–text pair datasets, containing RS images (RGB) from different regions, resolutions, and scene types. Due to large variations in image size and relatively low data quality, we apply simple preprocessing and filtering, resulting in a curated subset of about 1.04M images.
  • USA-1m (multispectral) Our self-constructed, urban-scale remote sensing dataset consisting of 8.77M four-channel, high-resolution ($1$ m) images, derived from the USDA National Agriculture Imagery Program (NAIP) [28]. The dataset covers major urban areas across 45 U.S. states, with a total footprint of $1.34339\times 10^{6}$ km2.
@@
In Stage II, we adopt five scene-classification datasets with distinct characteristics: NWPU-RESISC45 [7], UC Merced Land-Use [46], AID [40], WHU-RS19 [2], and RSD46-WHU [23]. Detailed information regarding their resolution, dataset size, number of classes, and data sources can be found in supplementary materials.
To comprehensively assess the impact of training data on RS generation models, we evaluate their generation capability directly on the test set and further use downstream-task metrics to evaluate the capability of the pretrained models, i.e., super-resolution (SR) and semantic image synthesis (SIS). The evaluation metrics include FID [12] and LPIPS [49]. Lower LPIPS and FID values indicate superior generation performance.
  • SECOND-SR We construct a super-resolution dataset based on the high-resolution SECOND semantic change detection benchmark [44]. Low-resolution (LR) images are generated by first downsampling and then upsampling the original high-resolution (HR) images. The dataset contains 4,662 paired LR–HR images, each with a spatial size of $512{\times}512$ pixels, covering six major land-cover classes.
  • OpenEarthMap [41] A global benchmark designed for semantic segmentation. It consists of 5,000 RS images across six continents at 0.25 to 0.5 m resolutions. Labels are annotated with eight land-cover categories.
@@ 4.2 Implementation Details
All models are implemented using the official DiT framework [25], which is a transformer-based architecture using its github repository [^1] https://github.com/facebookresearch/DiT. We adopt the DiT-XL/2 backbone for both generation pretraining and downstream tasks. All experiments are conducted with $256{\times}256$ input size, a global batch size of 256, and AdamW optimizer with a fixed learning rate of $1\times 10^{-4}$.
@@ 4.3 Comparison Results
We compare the proposed method with three groups of pruning strategies: basic pruning methods (e.g., Random, Moderate-DS [42]), data pruning methods for foundation models (e.g., Dinov2 [24], A2-MAE [48]), and a recent approach specifically developed for generation models ($\mathrm{Cluster}_{\text{nearest}}$ [6]). To examine their effectiveness across varying scene complexities, experiments are conducted on a global-scale dataset (Git-10M) and an urban-scale dataset (USA-1m). To ensure comparability, all models are trained for an equivalent number of iterations, 100K for Git-10M and 40K for USA-1m, and the best results during training are reported. Beyond generation evaluation on the test set, pretrained models are further assessed on two downstream tasks, i.e., SR and SIS, with the top fine-tuning results recorded.

### v2

@@ 1 Introduction
In recent years, generation models, especially diffusion models [Peebles2022DiT, Ho2020DDPM], have achieved remarkable progress in fields such as computer vision [Richard2021Binaural], medical imaging [Song2021ScoreMedical], and remote sensing (RS) [Dong2024BridgeSR]. Within the RS domain, generation models have been widely applied to data augmentation, image reconstruction, super-resolution, and high-resolution image synthesis, supporting practical applications in urban planning, land-use monitoring, and disaster response [Borana2023UrbanSustainability]. A powerful RS diffusion foundation model can provide a robust data and modeling backbone to further enhance these applications.
However, the effective training of RS diffusion foundation models critically depends on the quality and distribution of training data. The emergence of large-scale open-source datasets (e.g., Git-10M [Liu2025Text2Earth], RS5M [Zhang2024RS5M]) provides valuable resources, yet it also introduces several critical challenges, including image redundancy, low-quality samples (e.g., noise and cloud cover), class imbalance [Cheng2017NWPU], and scene homogeneity [Xia2017AID]. These issues not only hinder the training efficiency of RS foundation models but also limit the effectiveness of the resulting pretrained models on downstream tasks.
Recently, several studies in RS generation tasks have made preliminary explorations into data processing. For instance, RSDiff employs size cropping and noise augmentation [Sebaq2024RSDiff]. WHU-RS19 [Balestra2025WHURS19] removes cloud and open-ocean regions to reduce low information content. Super-resolution works [Huang2015Urban100] often prioritize urban areas. Nevertheless, these methods mainly rely on rule-based or simplistic strategies to eliminate redundant data, without considering the specific dependence of generation models on data distribution [Goodfellow2014GAN] or the inherent characteristics of RS imagery, such as diversity, heterogeneity, and class balance. Moreover, many data pruning methods developed in the computer vision domain rely on scoring mechanisms [Yang2024CLIPDataSelection] from supervised pretrained models [Pleiss2020AUCMargin] or retrieval of nearby images based on ImageNet benchmarks [Oquab2023DINOv2]. In contrast, a standardized labeled dataset for RS, comparable to ImageNet [Deng2009ImageNet] in computer vision and spanning multiple resolutions and modalities, is currently lacking. Overall, systematic research on data selection for RS diffusion foundation models remains scarce.
@@
The effectiveness of diffusion foundation models depends more on the quality and distribution of the training data than on dataset size alone. Recent studies [Briq2024DataPruning] have shown that diffusion models retain strong generation performance even when a large fraction (e.g., 90% of ImageNet samples) of the training data is removed, highlighting that a significant portion of the data contributes minimally. In the RS domain, raw data often contains substantial redundancy and noise, meaning that merely increasing the dataset size does not guarantee proportional performance gains. If generative models are trained on large volumes of data without appropriate selection, low-quality and redundant samples can not only slow convergence and increase computational and time costs, but also introduce distribution shifts that ultimately degrade model performance.
@@ 2.1 Diffusion foundation Models
Diffusion foundation models [Saharia2022Imagen] play a pivotal role in image generation, data synthesis, and image reconstruction. For instance, Stable Diffusion (SD) 1.5 [Rombach2022LDM] leverages latent-space representations together with a UNet backbone, serving as a lightweight generation foundation model that provides strong pretrained initialization for ControlNet [Zhang2023ControlNet] and related low-level vision tasks [Saharia2022Palette]. More recently, Transformer-based architectures such as DiT [Peebles2022DiT] have emerged, while models like SD3 [Esser2024RectifiedFlow] and Flux [BFL2024FluxRepo] offer substantially stronger generation capabilities and higher fidelity, further advancing the scalability and versatility of diffusion frameworks.
In the RS domain, diffusion foundation models have also started to gain attention. To address the unique properties of RS imagery, such as multi-spectral and multi-resolution observations, geospatial information, and global coverage, many works have explored pre-training or fine-tuning foundation models on RS datasets [Tang2024CRSDiff, Toker2024SatSynth, Xiao2023EDiffSR]. For example, DiffusionSat [Khanna2023DiffusionSat] introduced geolocation as conditioning information and fine-tuned Stable Diffusion 1.5 on multi-source RS data, supporting multiple RS generation tasks. Meanwhile, models such as SR3 [Saharia2021SR3] and CDM [Ho2021CDM], demonstrated strong super-resolution performance on natural images, providing a foundation for their adaptation to RS. Despite these advances, existing RS generation models typically aggregate multiple classification datasets or apply simplistic deduplication [Liu2024DiffusionSurvey], thereby overlooking the distributional requirements of generation foundation modeling as well as the inherent heterogeneity and diversity of RS imagery. Consequently, the field lacks systematic exploration of data pruning strategies specifically designed to address the characteristics of remote sensing data and the requirements of RS generative foundation models.
@@ 2.2 Data Pruning Methods
Training data are critical to constructing RS diffusion foundation models. However, existing studies [Liu2023RemoteCLIP, Sebaq2024RSDiff] in this field have adopted simplistic deduplication and preprocessing strategies, such as removing cloud and ocean regions [Balestra2025WHURS19] and prioritizing urban areas [Zhang2024A2MAE]. These approaches are insufficient to address the sensitivity of generation models to data distribution, heterogeneity, and class imbalance.
Current data pruning methods can be broadly categorized into three types. First, data-valuation methods assign an importance score to each sample and select samples accordingly. For example, MoSo [Tan2023MoSo] estimates the change in empirical risk when a sample is removed. Although such methods are generally efficient, their performance can be affected by group effects and may lack generalization in complex real-world settings. Second, distribution-based methods rely on the geometric structure of the dataset. For instance, Moderate-DS [moderate] selected samples near the median. CCS [Zheng2022CCS] balanced data distribution and sample importance during selection. Finally, optimization-based methods leverage optimization techniques to guide sample pruning, such as temporal dual-depth scoring [Zhang2024TDDS], gradient matching [Killamsetty2021GradMatch], scalable self-supervised pruning metrics [Sorscher2022DataPruning], influence functions [KohLiang2017Influence], and bilevel optimization [Borsos2020CoresetBilevel].
@@
Beyond explicit pruning algorithms, recent foundation models have also explored large-scale data pruning pipelines. For example, DINOv2 [Oquab2023DINOv2] constructs its curated LVD-142M dataset by performing PCA-hash and copy-detection based deduplication, followed by self-supervised retrieval based on ImageNet benchmarks. Similarly, V-JEPA 2 [assran2025vjepa2selfsupervisedvideo] adopts a retrieval and reweighting strategy to align large-scale video data with target distributions.
@@ 3.2 Stage I: Entropy-based Pruning
We measure the information value of each image by computing its global Shannon entropy [Shannon1948]. Specifically, for an image II, we compute its grayscale entropy $H(I)$ to capture the diversity of pixel intensities:
@@ 3.3 Stage II: Scene-Aware Clustering and Sampling
We employ a unified feature extractor $f(\cdot)$, namely Git-RSCLIP [Liu2025Text2Earth], with $\ell_{2}$ normalization to embed all images from standard datasets. Git-RSCLIP is specifically pretrained on large-scale remote sensing corpora and achieves substantially better performance than previous RS-oriented CLIP variants such as RemoteCLIP and GeoRSCLIP.
@@ 4.1 Datasets and evaluation
Moderate-DS [moderate] 44.07 92.40 0.3952 132.85 0.6222 199.24 124.86 0.5098 233.13 0.6730
$\mathrm{Cluster}_{\text{nearest}}$ [Briq2024DataPruning] 45.24 89.58 0.3898 133.18 0.6237 201.36 150.11 0.5870 256.40 0.6804
Dinov2 [Oquab2023DINOv2] 39.34 91.76 0.3930 137.55 0.6121 180.98 127.73 0.5472 238.80 0.6761
A2-MAE [Zhang2024A2MAE] 45.96 90.39 0.3938 130.64 0.6013 360.74 148.36 0.6028 256.39 0.6993
@@
To evaluate the effectiveness of our data pruning approach for RS generation tasks, we conduct experiments on both global-scale and urban-scale datasets. Specifically, we use two representative global-scale optical datasets (i.e., GIT-10M and RS5M) that were carefully curated for RS generative foundation models, as well as a large-scale urban-scale multispectral dataset (i.e., USA-1m) collected in this work based on U.S. urban boundary products [Li2020GUB]. The global datasets provide broad coverage and diversity, while the urban dataset offers finer spatial resolution and richer texture information. This design allows us to assess the effectiveness and generalizability of our data pruning strategy across datasets with varying resolutions, modalities, volumes, and sensors.
  • Git-10M [Liu2025Text2Earth] A global-scale dataset comprising 10.5M satellite images (RGB) with a spatial resolution of 0.5 to 128 m, which spans multiple continents and geographic regions, covering diverse land-cover types such as urban areas, croplands, forests, mountains, and deserts.
  • RS5M [Zhang2024RS5M] A global-scale dataset initially collected from 11 publicly available image–text pair datasets, containing RS images (RGB) from different regions, resolutions, and scene types. Due to large variations in image size and relatively low data quality, we apply simple preprocessing and filtering, resulting in a curated subset of about 1.04M images.
  • USA-1m (multispectral) Our self-constructed, urban-scale remote sensing dataset consisting of 8.77M four-channel, high-resolution ($1$ m) images, derived from the USDA National Agriculture Imagery Program (NAIP) [Robinson2019LargeScale]. The dataset covers major urban areas across 45 U.S. states, with a total footprint of $1.34339\times 10^{6}$ km2.
@@
In Stage II, we adopt five scene-classification datasets with distinct characteristics: NWPU-RESISC45 [Cheng2017NWPU], UC Merced Land-Use [Yang2010UCMerced], AID [Xia2017AID], WHU-RS19 [Balestra2025WHURS19], and RSD46-WHU [Long2017RSD46]. Detailed information regarding their resolution, dataset size, number of classes, and data sources can be found in supplementary materials.
To comprehensively assess the impact of training data on RS generation models, we evaluate their generation capability directly on the test set and further use downstream-task metrics to evaluate the capability of the pretrained models, i.e., super-resolution (SR) and semantic image synthesis (SIS). The evaluation metrics include FID [Heusel2017FID] and LPIPS [Zhang2018LPIPS]. Lower LPIPS and FID values indicate superior generation performance.
  • SECOND-SR We construct a super-resolution dataset based on the high-resolution SECOND semantic change detection benchmark [Yang2020SECOND]. Low-resolution (LR) images are generated by first downsampling and then upsampling the original high-resolution (HR) images. The dataset contains 4,662 paired LR–HR images, each with a spatial size of $512{\times}512$ pixels, covering six major land-cover classes.
  • OpenEarthMap [Xia2023OpenEarthMap] A global benchmark designed for semantic segmentation. It consists of 5,000 RS images across six continents at 0.25 to 0.5 m resolutions. Labels are annotated with eight land-cover categories.
@@ 4.2 Implementation Details
All models are implemented using the official DiT framework [Peebles2022DiT], which is a transformer-based architecture using its github repository [^1] https://github.com/facebookresearch/DiT. We adopt the DiT-XL/2 backbone for both generation pretraining and downstream tasks. All experiments are conducted with $256{\times}256$ input size, a global batch size of 256, and AdamW optimizer with a fixed learning rate of $1\times 10^{-4}$.
@@ 4.3 Comparison Results
We compare the proposed method with three groups of pruning strategies: basic pruning methods (e.g., Random, Moderate-DS [moderate]), data pruning methods for foundation models (e.g., Dinov2 [Oquab2023DINOv2], A2-MAE [Zhang2024A2MAE]), and a recent approach specifically developed for generation models ($\mathrm{Cluster}_{\text{nearest}}$ [Briq2024DataPruning]). To examine their effectiveness across varying scene complexities, experiments are conducted on a global-scale dataset (Git-10M) and an urban-scale dataset (USA-1m). To ensure comparability, all models are trained for an equivalent number of iterations, 100K for Git-10M and 40K for USA-1m, and the best results during training are reported. Beyond generation evaluation on the test set, pretrained models are further assessed on two downstream tasks, i.e., SR and SIS, with the top fine-tuning results recorded.